OpenStudy (anonymous):

Is this true: 'The only way you can get from one perfect square, q^2, to another, p^2, by multiplying by a constant [i.e. p^2=a^2 q^2] is if q IS that factor [i.e. q^2=a^2]'?

hartnn (hartnn):

'the first square IS that factor' what is 'that' here? factor of whose ? \(b^2=ma^2\)

OpenStudy (anonymous):

\[q^2=49p^2\]Must \[q=49\]?

OpenStudy (anonymous):

I've edited it. Is that still ambiguous?

hartnn (hartnn):

then i don't think its always true.... q equal 49, is not necessary

OpenStudy (anonymous):

Of course apart from the trivial case where p=1

OpenStudy (anonymous):

Counterexample: 6^2= 4 * 2^2 Damn, it's wrong!

hartnn (hartnn):

yea, its false....

OpenStudy (anonymous):

This popped up in http://www.komal.hu/verseny/feladat.cgi?a=honap&h=201210&t=mat&l=en if anyone's interested...